OpenStudy (anonymous):

Help with these, please? 1. simplify 5x^2+9x-2/x^2+12x+20 times x^2+17x+70/15 x-3 2. simplify 25w^6/10w^3 divided by 30w^2/5w 3. simplify 8x-56/x^2-49 divided by x-6/x^2+11x+28 4. x^2+x-12/x^2-x-20 divided by 3x^2-24x+45/12x^2-48x-60

OpenStudy (mertsj):

\[\frac{5x^2+9x-2}{x^2+12x+20}\times\frac{x^2+17x+70}{15x-3}\]

OpenStudy (mertsj):

Factor all expressions and divide out the common factors.